### v2

# DriveGen3D: Boosting Feed-Forward Driving Scene Generation with Efficient Video Diffusion

Abstract
We present DriveGen3D, a novel framework for generating high-quality and highly controllable dynamic 3D driving scenes that addresses critical limitations in existing methodologies. Current approaches to driving scene synthesis either suffer from prohibitive computational demands for extended temporal generation, focus exclusively on prolonged video synthesis without 3D representation, or restrict themselves to static single-scene reconstruction. Our work bridges this methodological gap by integrating accelerated long-term video generation with large-scale dynamic scene reconstruction through multimodal conditional control. DriveGen3D introduces a unified pipeline consisting of two specialized components: FastDrive-DiT, an efficient video diffusion transformer for high-resolution, temporally coherent video synthesis under text and Bird’s-Eye-View (BEV) layout guidance; and FastRecon3D, a feed-forward module that rapidly builds 3D Gaussian representations across time, ensuring spatial-temporal consistency. DriveGen3D enable the generation of long driving videos (up to $800\times 424$ at $12$ FPS) and corresponding 3D scenes, achieving state-of-the-art results while maintaining efficiency.
I Introduction
The synthesis of 3D dynamic driving environments has emerged as a key research frontier in autonomous systems, driven by its wide-ranging applications in simulation, perception, and planning. While recent advances in video generation [1] and 3D scene reconstruction [2, 3] have made substantial progress, an integrated and efficient framework that unifies long-horizon video synthesis and large-scale 3D scene reconstruction under multi-modal control is still lacking.
Existing methodologies typically address either temporal coherence in video generation or spatial fidelity in scene reconstruction, but not both—often requiring high computational cost or suffering from limited scalability. For example, state-of-the-art diffusion-based models like MagicDriveDiT [1] can produce high-resolution driving sequences but require up to 30 minutes to generate a single 20s video, making them impractical for real-time use.
From the reconstruction perspective, optimization-based approaches [2, 4] are similarly time-consuming, often needing 30 minutes per scene. While recent feed-forward methods [5, 6, 3, 7, 8] have reduced reconstruction time to seconds, they remain limited in scale and rarely integrate with dynamic video generation pipelines.
To bridge this gap, we propose DriveGen3D, an efficient and unified framework that integrates two specialized modules: FastDrive-DiT, an accelerated video diffusion transformer for high-resolution driving video generation, and FastRecon3D, a feed-forward reconstruction pipeline that builds dynamic 3D scenes from multi-view video frames in real-time.
To address the latency in video generation, we analyze the diffusion process and identify that applying caching strategies [9] exclusively to the conditional branch maximizes speedup without compromising layout control. Furthermore, we addressed the computational bottleneck of the attention layers. By profiling the attention maps, we apply precise 8-bit quantization [10, 11] specifically to the cross-view attention modules, reducing inference time by over $2\times$ while preserving visual fidelity.
FastRecon3D enables rapid driving scene reconstruction. Unlike prior methods that treat frames independently [5, 6] or optimize each scene individually [2], we propose a recursive, feed-forward and temporal-aware approach. By employing cross-attention to fuse features from past and future contexts, our model reconstructs 3D primitives that are both spatially accurate and temporally coherent. This effectively mitigates motion blur in dynamic elements and ensures consistent geometry across the synthesized sequence.
We conduct extensive experiments on the nuScenes dataset [12] demonstrating that DriveGen3D outperforms state-of-the-art baselines in both video quality and 3D reconstruction accuracy. Specifically, Our framework achieves a PSNR of 23.25, an SSIM of 0.820, and an LPIPS of 0.176 on generated inputs, significantly outperforming previous state-of-the-art methods. Moreover, DriveGen3D reduces the end-to-end generation time to under 6 minutes, including video synthesis and 3D scene reconstruction, marking a substantial improvement in efficiency. Our contributions can be summarized as follows:
Figure 1: Overview of DriveGen3D. (a) Given textual and BEV layout conditions, our model first employs an accelerated Video Diffusion Transformer to synthesize a long driving video. (b) Next, a per-frame 3D Gaussian Splatting representation is utilized to construct entire scene from the generated video frames.
  • We introduce DriveGen3D, a unified framework for efficient 3D driving scene generation that combines video synthesis and feed-forward 3D reconstruction under multimodal control via text and BEV layout.
  • We propose FastDrive-DiT, a high-performance video diffusion transformer equipped with step caching and quantized attention, achieving over $2\times$ inference speedup while maintaining visual fidelity. We also develop FastRecon3D, a feed-forward 3D scene reconstruction module based on temporal-aware Gaussian splatting, enabling fast and scalable scene modeling from generated videos.
  • DriveGen3D achieves high-quality video generation and complete 3D reconstruction for large-scale street scenes within 6 minutes, significantly outperforming the state-of-the-art methods in terms of photorealism, scalability and efficiency.
II Related Work
II-A Video generation for driving scene
Recent advancements in street view generation and autonomous driving scene synthesis have significantly improved the fidelity and controllability of synthetic data. MagicDrive [13] introduces a framework for street view generation with diverse 3D geometry controls, such as camera poses and 3D bounding boxes, enhancing tasks like BEV segmentation and 3D object detection through cross-view consistency. MagicDriveDiT [1] extends this by addressing high-resolution, long video generation for autonomous driving, leveraging flow matching and spatial-temporal conditional encoding to achieve superior scalability and control. DrivingDiffusion [14] synthesizes multi-view driving videos with precise layout control, ensuring cross-view and cross-frame consistency through a spatial-temporal diffusion framework. UniScene [15] unifies the generation of semantic occupancy, video, and LiDAR data, employing a progressive generation process to reduce complexity and improve downstream task performance.Together, these works advance the field of autonomous driving by providing scalable, controllable, and high-fidelity synthetic data generation frameworks.
II-B Reconstruction for driving scene
The reconstruction of dynamic driving scenes [2] has emerged as a critical task in autonomous systems and immersive environment modeling. Contemporary approaches predominantly leverage Gaussian splatting-based representations due to their inherent balance between rendering efficiency and geometric expressiveness. Early methodologies in this domain adopted optimization-based paradigms, exemplified by works such as StreetGaussian [2] and DrivingGaussian [4]. These frameworks optimize scene representations per instance for specific street segments (typically under 100 meters in scale) through iterative refinement processes spanning approximately 30 minutes. Recent advancements have shifted toward feed-forward architectures [5, 6, 16, 17, 18, 19, 20] to enable rapid 3D reconstruction. Methods like PixelSplat [5] and MVSplat [6] employ large pretrained networks to directly infer Gaussian parameters from multi-view inputs, reducing reconstruction time from minutes to seconds. Though these approaches demonstrate generalizability across scenes, they often sacrifice reconstruction fidelity in geometrically complex regions or under sparse observational constraints. Parallel innovations address the temporal dimension of driving scenes: InfiniCube [21] extends the Scube [22] framework to 3D street generation, disentangling dynamic vehicles from static backgrounds via hybrid voxel-video control mechanisms. Drive3R [8] adapts the architecture of Spann3R [23] for per-frame 3D scene reconstruction through temporal consistency priors. DrivingRecon [7] mimics StreetGaussians [2] but replaces iterative optimization with feed-forward prediction. DrivingForward [3] enhances sparse-view reconstruction robustness by jointly learning pose estimation and depth prediction modules within its network architecture. These advancements collectively highlight two persistent limitations:
  • Existing feed-forward 3D reconstruction methods operate at constrained spatial resolutions (typically $\leq 512\times 512$)
  • Prior works on generative models for conditional scene synthesis required substantial computational resources and complex framework integration, which hindered their widespread adoption.
III Method
III-A Overview
DriveGen3D is an integrated 3D driving scene generation framework composed of two key components as illustrated in Figure 1. The pipeline begins with FastDrive-DiT, which synthesizes high-resolution, temporally coherent driving videos under conditional guidance. These generated videos are then passed to FastRecon3D, which reconstructs dynamic 3D scenes in a feed-forward manner using temporal-aware Gaussian splatting. Together, these modules enable rapid and scalable 3D scene generation suitable for real-time simulation and autonomous driving applications.
The architecture of our efficient video diffusion module, dubbed FastDrive-DiT, integrates two targeted acceleration strategies to address the high inference cost of driving video generation, leading to a speedup of more than $2\times$. FastRecon3D proposes a temporally context-visible fusion module based on the static feed-forward framework, resulting in better temporally consistent per-frame Gaussians.
III-B FastDrive-DiT
Generating 3D driving scenes, especially in the autonomous driving domain, is notoriously time-consuming due to the multi-view nature of the data. The video generation step is particularly costly because of the underlying diffusion process. For instance, MagicDriveDiT [1] can take up to 30 minutes to produce a video of driving scene. To address this inefficiency, we propose FastDrive-DiT, an enhanced and lightweight video diffusion model built on MagicDriveDiT with targeted acceleration strategies.
Diffusion steps acceraleration. To accelerate the generation of long and high-resolution videos in MagicDriveDiT, we integrate TeaCache[9], a training-free caching approach that enhances the inference speed of diffusion models. TeaCache leverages timestep embeddings to modulate noisy inputs, approximating the fluctuating differences in model outputs across timesteps. By introducing a rescaling strategy, it refines these differences, enabling efficient output caching with minimal computational overhead. Experiments show that TeaCache achieves up to $4.41\times$ acceleration while maintaining visual quality, making it an effective solution for balancing speed and performance in video generation.
In the original TeaCache[9] implementation, coefficients are computed from both the conditional and unconditional branches of the diffusion model. However, in our adaptation for MagicDriveDiT, we calculate coefficients exclusively from the conditional branch. This optimization not only reduces the forwarding time of TeaCache but also ensures no obvious degradation in performance, further enhancing the efficiency of the video generation process.
Quantized DiT. Considering the high cost of the attention part in the video DiT, we use SageAttention[10, 11] for quantization during the inference procedure.
[…]
We carefully analyze the inference profiling and visualize $Q\times K\times V$ of different attention blocks, pointing out potential improvement techniques. In this paper, applying sageattention to the transformer blocks saves an additional 30 seconds of inference time with nearly no performance degradation. Experiment details are in section 6.
Figure 2: Visualization of the multi-view reconstructed video from a generated 3D scene.
III-C FastRecon3D
While the aforementioned methods enable realistic driving scene generation, applications like simulation require complete 3D scene models. To enable rapid novel scene synthesis, we introduce FastRecon3D, a feed-forward reconstruction module that avoids costly optimization while preserving quality. Built upon temporal-aware Gaussian Splatting, FastRecon3D reconstructs the 3D geometry of each frame 3D geometry by leveraging both past and future contexts, enabling fast and consistent 3D scene generation.
Per-Frame 3D Representation. Existing methods like Street Gaussians [2] and Infinicube [21] attempt dynamic scene reconstruction through object-scene separation but suffer from two critical limitations: 1) Dynamic elements in static backgrounds (e.g., pedestrians, cyclists) suffer motion blur artifacts 2) Computational constraints prevent large-scale scene reconstruction. Following DrivingForward [3], we propose a temporal-aware Gaussian Splatting formulation that reconstructs per-frame 3D Gaussian models while maintaining 3D consistency.
In our approach, each time step is represented by a set of 3D Gaussian primitives:
$G_{i}^{t}=\{\mathcal{G}_{i}^{t}\}_{i=1}^{N_{t}}=\{(\mathbf{\mu}_{i},\mathbf{\Sigma}_{i},\mathbf{\alpha}_{i},\mathbf{c}_{i})\}_{i=1}^{N_{t}}.$ (1)
Recursively reconstruction from videos. We employ the decoder of DrivingForward [3] as the Gaussian Network to estimate Gaussian parameters, and use a pre-trained image encoder [25] to extract features.
After the frame generation of the $t+\Delta$ timestep, we extract all frames from time steps $t-\Delta$ to $t+\Delta$ to reconstruct the scene at time $t$. Our model first extracts features $\{F_{i}^{t}\}_{i=1}^{N_{t}}$ from each frame. For each timestep $t$, given multi-view features $\{F_{i}^{t}\}_{i=1}^{N_{t}}$ and temporal neighbors $\{F_{i}^{t\pm\Delta}\}_{i=1}^{N_{t}}$, we use cross-attention to fuse information from other frames into $\{F_{i}^{t}\}_{i=1}^{N_{t}}$:
$\{{F^{\prime}}_{i}^{t}\}_{i=1}^{N_{t}}=\text{Attention}(Q,K,V),$ (2)
$Q\leftarrow\{F_{i}^{t}\}_{i=1}^{N_{t}},\quad K,V\leftarrow\{F_{i}^{t\pm\Delta}\}_{i=1}^{N_{t}}.$ (3)
TABLE I: Comparison of our method against prior feed-forward and optimization-based methods. The last two rows show novel view rendering performance with either GT or generated video input.
Type Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Static MVSplat [6] 22.83 0.629 0.317
pixelSplat [5] 25.00 0.727 0.298
Dynamic UniPad [26] 16.45 0.375 -
SelfOcc [27] 18.22 0.464 -
EmerNeRF [28] 20.95 0.585 -
DistillNeRF [29] 20.78 0.590 -
DrivingForward [3] 21.69 0.742 0.226
Ours (w/ GT images) 23.87 0.814 0.169
Ours (w/ GEN images) 23.25 0.820 0.176
Then the network predicts Gaussian parameters and save it as a 3D model:
$\mathcal{G}^{t}=\{\mu^{t},\Sigma^{t},\alpha^{t},c^{t}\}=\Psi\Bigl(\{{F^{\prime}}_{i}^{t-\Delta},\,{F^{\prime}}_{i}^{t},\,{F^{\prime}}_{i}^{t+\Delta}\}_{i=1}^{N_{t}}\Bigr).$ (4)
By leveraging both past and future context, this recursive reconstruction method effectively captures dynamic scene elements while maintaining high spatial fidelity. As a result, our approach produces complete 3D models in a matter of seconds, meeting the demanding requirements of simulation and other real-time applications without compromising on quality.
IV Experiments
IV-A Experimental Setup
Dataset. The training dataset is obtained from the nuScenes dataset [12]. It consists of 700 training videos and 150 validation videos. For the 3D scene reconstruction model, we split the dataset into 20,000 short sequences for training.
Figure 3: Typical examples of the data distribution of tensors in different attention blocks of MagicDriveDiT.
(a) All
(b) Condition
(c) Uncondition
Figure 4: Visualization of input differences and output differences in consecutive timesteps. We plot all, condition and uncondition separately.
Evaluation Metrics. For video-generation stage, we evaluate both the realism and controllability in street-view video generation. For video generation, we adhere to the benchmarks from [30]. To measure video quality, we use the Frechet Video Distance (FVD). Regarding controllability, we employ mAP from 3D object detection and mIoU from BEV segmentation. For both of these tasks, we utilize BEVFormer [31], which is a video-based perception model. We generate corresponding videos for the annotations in the validation set. Then, we apply the aforementioned metrics and use perception models pre-trained on real-world data for evaluation.
For 3D scene reconstruction stage, we adopt the view synthesis task to assess reconstruction quality, following the evaluation protocol established in Driving Forward [3]. Given sequential video frames from timestamps $t-\Delta$ and $t+\Delta$ (In our implementation, $\Delta=2$) as input, our model synthesizes the intermediate frame at timestamp $t$ for quantitative comparison against ground truth. We report Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM), and Learned Perceptual Image Patch Similarity (LPIPS) [32] in Table I.
TABLE II: Time cost of different attention blocks of MagicDriveDiT during inference.
Total spatial temporal cross-view cross other
615 s 104 s 82 s 163 s 62 s 204 s
TABLE III: Acceralerating the inference process of MagicDriveDiT. 17f and 233f denote the frames count of generated videos.
[…]
MagicDriveDiT 111.58 17.10 21.92 211s/615 s
Ours (w/o Quant) 125.70 16.60 21.27 64 s/309 s
Ours 125.88 16.72 21.24 58 s/278 s
To benchmark performance, we compare our DriveGen3D reconstruction model against the Driving Forward baseline (evaluate on ground truth images) across same validation frames and synthesized video outputs. This dual evaluation strategy disentangles reconstruction capability from generation artifacts, providing comprehensive insights into geometric and appearance recovery accuracy.
Implementation Details. We train the model with a resolution of $800\times 424$. Inference is performed on NVIDIA H20 GPUs. When assessing the time cost of our proposed method, the baselines for the video generation model are MagicDriveDiT [1](17f) and MagicDriveDiT [1](233f). The 3D scene reconstruction stage is trained on a single NVIDIA H20 GPU for 2 days.
IV-B Experimental Results
IV-B1 Novel View Synthesis
Table I provides a comprehensive comparison of DriveGen3D against both optimization-based dynamic methods and feed-forward reconstruction methods. Notably, when using generated images instead of ground truth, DriveGen3D maintains competitive reconstruction quality with a PSNR of 22.84 and achieves the highest SSIM of 0.811, demonstrating strong temporal coherence and structure preservation in generated scenes. This suggests that despite operating on synthetic inputs, DriveGen3D produces reliable and structurally consistent 3D reconstructions, validating the effectiveness of its end-to-end pipeline.
IV-B2 Experimental Analysis and Ablation Study
We construct the following experiments to demonstrate the effectiveness of FastDrive-DiT.
Diffusion steps acceraleration. Firstly, we visualize the input differences and output differences in consecutive timesteps in Figure 4. It is observed that the default configurations of TeaCache [9], i.e. all, exhibits a U-shape for model output, with downward tend initially, nearly constant for the middle and upward until the end. The same is for the uncondition branch. The condition branch exhibits a slightly different phenomenon, with the start of model output much smaller. We also apply simple polynomial fitting to fit a relationship between model input and output and the use these cofficients to predict model output according to the input.
Figure 5: Comparison of video generation for MagicDriveDiT, Diffusion steps acceleration and Quantized DiT.
Figure 6: Comparison of GT videos and generated video for reconstruction.
As shown in the black lines in Figure 4, the quality of fitting is best for the condition branch, while all and uncondition do not fit well for the start. We attribute this to the different input and output relationships in the start of diffusion process. So we finally only apply TeaCache to the condition branch of MagicDriveDiT. As shown in Table III, equiping MagicDriveDiT with TeaCache for condition branch has a speedup of nearly $3\times$ and $2\times$ for 17 frames and 233 frames. The perception metrics, mAP and mIoU only shows slight decrease.
Quantized DiT. Secondly, we profile the time cost of MagicDriveDiT. As shown in Table  II, cross-view attention is identified as the most computationally costly process in MagicDriveDiT. We further analyze different attention components in MagicDriveDiT and plot the distribution of $Q,K,V$, as depicted in Figure 3. Notably, $Q$ and $K$ exhibit a trend similar to that in Figure 4 of SageAttention2 [11]. An interesting phenomenon is observed: the numeric range of $V$ follows the order: spatial >> temporal >> cross-view. Specifically, the cross-view range is 10 times smaller than the spatial range and 5 times smaller than the temporal range. Given the principle of quantization, a smaller range is more conducive to quantization. Therefore, a promising solution is not only to apply SageAttention [10] to cross-view attention but also substitute the high-precision quantization method for V with a lower-resolution one. Concretely, for $V$, the FP8, E4M3 data type can be replaced with FP8, E5M2 data type and INT8. We leave the latter for future attempt. In this paper, applying sageattention to the transformer blocks saves an additional 30 seconds (233 frames) of inference time after diffusion steps acceraleration while maintaining performance.
IV-B3 Visualization and Comparison
We provide visual results of DriveGen3D alongside ablation experiments.
Visualization. In Figure 5, we compare the generated videos of baseline model, baseline model with diffusion steps acceraleration and further quantization with SageAttention. With the proposed techniques, no obvious change is observed. It can still generate video longer than 15 s with a much faster speed. This highlights the efficiency of DriveGen3D.
Full pipeline results. In Figure 2, we show a typical output result of our framework. DriveGen3D can generate and reconstruct street scenes for more than 20s, 12FPS.
Comparison with raw videos. By default, our reconstruction stage takes the generated videos as input. We compare with DrivingForward to validate the difference of raw videos and generated videos. As evidenced in Table I, DriveGen3D maintains fidelity in video generation that is essentially comparable to existing methodologies with GT images, exhibiting only a marginal degradation. Qualitative analysis is shown in Figure 6. In the front view image of reconstruction videos, the spatial location of generated bus is different from gt videos, we check the 3D Box layout and find the generation videos does not match well. In the view of back left, the generation images is blurry, underscoring the needs of optimizing video generation model for reconstruction. Moreover, both GT videos and generation videos display drawbacks of flickering pole, pointing out the flaw of second stage reconstruction model.
V Conclusion
We present DriveGen3D, an efficient framework for synthesizing long-duration driving videos from textual descriptions and BEV layouts, and generating high-quality large-scale dynamic scenes. Our approach marks a significant advancement in world modeling by bypassing conventional voxel-based generation paradigms [21] through a novel integration of longitudinal video generation and scene reconstruction modules. This architecture enables faithful reproduction of real-world driving scenarios and thus paves the way for novel applications in autonomous vehicle simulation and dynamic world modeling.